You are studying a population of 5000 winged Fargles. There are 3200 green Fargles and 1800 purple Fargles. Calculate p and q for the population
1
Expert's answer
2018-01-23T15:18:07-0500
If green is dominant:

q= sqrt (1800/5000)=sqrt(0.36)=0.6
p=1 - 0.6=0.4

If purple is dominant:
q= sqrt (3200/5000)=sqrt(0.64)=0.8
p=1 - 0.8=0.2